OpenStudy (anonymous):

Write the solution as either the union or intersection of two sets. 2│x-7│>28

OpenStudy (anonymous):

hello, so we can divide both sides by 2: |x-7| > 14 now we will check for what x's the next equation holds: |x-7| = 14 now we have two options : 1. x-7 = 14 or 2. x-7 = -14 so 1. x = 21 2. x = -7 now we have found the values for |x-7|=14 but we want |x-7|>14 so we sketch:

OpenStudy (anonymous):



OpenStudy (anonymous):

for example, for the x<-7 part we can choose x = -10 plug it into the inequality |-10-7| = -17 and it is >14 so it is good for us. for the -7<x<21 we can choose x =0 |0-7| = 7 and it is < 14 so it is not good for us either. we left with x>21 we choose 22 |22 - 7| = 15 is it >14 and it is good for us so the solution is x>21 or x<-7 ( union)